An object has a weight of 690 Newton on earth, specify (a) its mass in kilograms, (b) its weight in pound-force, and (c) its mass in slugs. If the object is on other celestial objects, where the gravitational acceleration is g = 5.50 ft/s2 , determine (d) its mass in kilograms, and (e) its weight in pounds-force.


1
Expert's answer
2021-03-26T11:37:25-0400

(a) Find its mass in kilograms:


"m=\frac Wg=70.4\text{ kg}."

(b) Find its weight in pound-force. 1 newton is 0.224809 pound-force, so,


"W'=0.2248W=155.1\text{ lbf}."

(c) Find its mass in slugs. 1 kg is 0.06852 slugs, so, 70.4 kg is 4.796 slug.



If the object is on other celestial objects, where the gravitational acceleration is g = 5.50 ft/s2,

(d) its mass in kilograms will not change and will remain 70.4 kg.

(e) Find its weight in pounds-force. We know that acceleration due to gravity on earth is 32.17 ft/s2. So, its weight on the celestial body will be


"W''=W'\frac{5.50}{32.17}=155.1·\frac{5.50}{32.17}=26.5\text{ lbf}."